OpenStudy (anonymous):

2. The strength of an electric field at 5.0 cm from a point charge is 100.0 N/C. What is the magnitude of the source charge?

OpenStudy (anonymous):

Look for a formula that uses these givens. D = Electric field = 100 N/C d = 5 cm but newtons are a kg m s unit so you have to change the distance to meters. d = 5 cm * [ 1 m / 100 cm ] = 0.05 meters k = 9 * 10^9 E = k q/(d^2 you are looking for q Finally, substitute. 100 = 9*10^9 q/(0.05)^2 100*0.0025 = 9*10^9 q 100*0.0025/9*10^9 = q q = 2.8*10^-11

OpenStudy (anonymous):

For short q=Er^2/k;k=1/4*pi*epsilon q=100*(0.05)^2/8.9x10^9=2.808x10^-11 coulombs